# Xingjian Wu, Jianxin Jin, Wanghui Qiu, Peng Chen, Yang Shu, Bin Yang, Chenjuan Guo🖂

### v4

@@ Abstract
Cross-domain generalization is very important in Time Series Forecasting because similar historical information may lead to distinct future trends due to the domain-specific characteristics. Recent works focus on building unimodal time series foundation models and end-to-end multimodal supervised models. Since domain-specific knowledge is often contained in modalities like texts, the former lacks the explicit utilization of them, thus hindering the performance. The latter is tailored for end-to-end scenarios and does not support zero-shot inference for cross-domain scenarios. In this work, we introduce Aurora, a Multimodal Time Series Foundation Model, which supports multimodal inputs and zero-shot inference. Pretrained on Cross-domain Multimodal Time Series Corpus, Aurora can adaptively extract and focus on key domain knowledge contained in corresponding text or image modalities, thus possessing strong cross-domain generalization capability. Through tokenization, encoding, and distillation, Aurora can extract multimodal domain knowledge as guidance and then utilizes a Modality-Guided Multi-head Self-Attention to inject them into the modeling of temporal representations. In the decoding phase, the multimodal representations are used to generate the conditions and prototypes of future tokens, contributing to a novel Prototype-Guided Flow Matching for generative probabilistic forecasting. Comprehensive experiments on 5 well-recognized benchmarks, including TimeMMD, TSFM-Bench, ProbTS, TFB, and EPF, demonstrate the consistent state-of-the-art performance of Aurora on both unimodal and multimodal scenarios.
@@ 1 Introduction
Time series forecasting has gained sustained attention for decades of years due to its significant values in multiple domains, including economy, transportation, meteorology, and public health. In recent years, the key pivot comes with the surge of deep learning, which brings the boom of merticulously-designed deep forecasting models (Cirstea et al., 2022; Nie et al., 2023; Qiu et al., 2025d; Wu et al., 2025d). Through learning the inherent dynamics within the raw data, deep learning models can outperform classic statistical methods (Box & Pierce, 1970; Mei et al., 2014) and obey the scaling law (Shi et al., 2024a; Yao et al., ). Due to the success, it also brings the most commonly-used forecasting paradigm, which utilizes the past information to infer how the series goes in the coming horizon. Although this paradigm contributes to impressive performance under the domain-specific scenarios, its effectiveness is uncertain when facing cross-domain inference, where similar historical information may lead to different futures due to domain differences.
@@
2) utilizing cross-modality information in training end-to-end multimodal supervised models, which effectively integrates domain knowledge in forecasting. For time series foundation models, the cross-domain generalization capabilities mainly come from the sensitivity to subtle differences in historical information from different domains. Some of them (Shi et al., 2024b; Liu et al., 2025e) are pretrained on trillion-scale corpus with hevay backbones, thus possessing certain cross-domain adaption capabilities. Others (Wang et al., 2025c; b; Wu et al., 2025a) have specific structures, which excels at capturing cross-domain features. However, their capabilites come from single time modality and lack explict domain knowledge guidance, thus hindering the performance. For end-to-end multimodal supervised models (Jin et al., 2024; Liu et al., 2025a), though they consider the multimodal knowledge to enhance the domain-specific forecasting, they lack the ability to support zero-shot forecasting in cross-domain scenarios. In our view, the aurora of next-generation time series foundation model lies in pretraining a cross-modality model on cross-domain time series corpus, which can utilize the domain knowledge within modalities and serve as a versatile out-of-the-box forecaster in complex scenarios.
See Figure 1, we propose Aurora by pionnering the exploration of multimodal time series foundation model. Specifically, we pretrain Aurora on Cross-Domain Multimodal Time Series Corpus, with time series data and sample-wise domain-specific text descriptions. Since previous works (Chen et al., 2024a; Yu et al., 2025) point out the endogenous images of time series contain additional geometric information, we also consider them into cross-modality learning. Considering the model architecture, Aurora adopts a novel cross-modality Encoder. Taking pretrained Bert (Devlin et al., 2019) and ViT (Liu et al., 2021) as modality encoders, Aurora then adopts token distillation to extract the key information in different modalities. To effectively model the cross-modality interaction, we propose a novel Modality-Guided Self-Attention Mechanism to utilize the external domain knowledge to adjust the attention of internal information within the time series data to obtain temporal features, and then fuse them with text and image features.
@@ 2.1 Time Series Forecasting
Time Series Forecasting is vital in decision-making and has fascinated people for decades of years, which facilitates the emergence of a series of works. In recent years, deep-learning models are widely studied, among them, Autoformer (Wu et al., 2021), Triformer (Cirstea et al., 2022), TimesNet (Wu et al., 2023), Pathformer (Chen et al., 2024b), PatchTST (Nie et al., 2023), Dlinear (Zeng et al., 2023), FiTS (Xu et al., 2024), SparseTSF (Lin et al., 2024a), PDF (Dai et al., 2024a), DUET (Qiu et al., 2025d), and TimeMixer++ (Wang et al., ), continuously advancing the state-of-the-arts. However, though they possess the capabilities to extract the inherent dynamics in raw time series data, they only adapt to unimodal end-to-end forecasting scenarios, and often fall short in multimodal forecasting scenarios where the domain knowledge is widely contained in the text modality.
Recently, some works are proposed to explore the multimodal end-to-end supervised models. In summary, they utilize Large Language Models’ strong reasoning capabilities to integrate textual domain knowledge to prompt temporal modeling. Among them, Unitime (Liu et al., 2024c), Time-LLM (Jin et al., 2024) and CALF (Liu et al., 2025a) utilize the endogenous textual descriptions as prompts, GTP4MTS (Jia et al., 2024), TATS (Li et al., 2025d) and TimeMMD (Liu et al., 2024b) supports exogenous textual domain knowledge. However, they do not possess generalization capabilities in zero-shot scenarios.
@@ 2.2 Time Series Foundation Models
To support cross-domain generalization, unimodal Time Series Foundation Models are widely studied. The majority of them adopt Tranformer-based architectures, which are pretrained on time series corpus of billion- or trillion- scale to obtain the strong generalization capabilities. Among them, Sundial (Liu et al., 2025e), VisionTS (Chen et al., 2024a), ROSE (Wang et al., 2025c), Time-MoE (Shi et al., 2024b), MOIRAI (Woo et al., 2024), TTM (Ekambaram et al., 2024), Chronos (Ansari et al., ), UniTS (Gao et al., 2024), Timer (Liu et al., 2024e), and TimesFM (Das et al., 2024) demonstrate strong zero-shot forecasting performance on unimodal tasks, even outperforming those full-shot supervised models in many cases. Considering the forecasting paradigm, Sundial, MOIRAI, Chronos, and Lag-Llama (Rasul et al., 2023) also support probabilistic forecasting, which provides additional robustness and versatility for decision-making. Despite their endeavors to enhance cross-domain generalization capabilities, when historical series exhibit similarities, the forecasts they generate remain static. This lack of adaptability renders them unable to accommodate the diverse and changing real-world domains.
@@ 3.1.3 Multimodal Alignment
where VAttn and TAttn are unnormalized attention scores, separately denoting the correlations between time modality and image or text modality. $\text{Corr}\in\mathbb{R}^{n^{time}\times n^{time}}$ denotes the inherent temporal correlations bridged through the cross-modality correlations. We also introduce $W\in\mathbb{R}^{K^{image}\times K^{text}}$ as a learnable metric (Qiu et al., 2025d) to further tune the semantic distances. This process can help bridge the correlations between time series tokens through multimodal information, capable of leveraging domain knowledege to guide the modeling of temporal dynamics. We then inject Corr into the temporal encoding process:
@@ 3.2.2 Prototype-Guided Flow Matching
where $\mathcal{D}$ denotes the weights of prototypes, we then generate the new prototypes through: $\tilde{\mathcal{P}}=\mathcal{D}\cdot\mathcal{P}\in\mathbb{R}^{F\times p^{time}}$, where the generated prototype $\tilde{\mathcal{P}}$ contains the approximate future periodicities and trends. As shown in Figure 3, the motivation of Flow Matching is to fit the velocity field between the initial prototype $y_{i}^{(0)}=\tilde{\mathcal{P}}_{i}+\epsilon_{i}$ and the target horizon $y_{i}^{(1)}=\mathrm{y}_{i}$, where $\mathrm{y}_{i}\in\mathbb{R}^{p^{time}}$ is the groundtruth of the $i$-th future token, and $\epsilon_{i}\sim\mathcal{N}(0,\mathbf{I})$ is used to increase the diversity during training. We design the Flow-Matching Network with an MLP structure and utilize the AdaLN (Peebles & Xie, 2023) to integrate the multimodal conditions $h_{i}=\mathrm{X}^{cond}_{i}$. We adopt the conditional optimal-transport path, which is energy-optimal and contributes to a uniform velocity field. And the function of Flow-Matching Network $v^{\theta}_{t}$ is to predict the velocity based on the current position $y_{i}^{(t)}$ and condition $h_{i}$. To achieve this, the token-wise optimization objective $\mathcal{L}$ is designed as:
@@
In the inference phase, the sampling process is a discretized integration process–see Algorithm 3.2.2. The Gaussian noise $\epsilon_{i}\sim\mathcal{N}(0,\mathbf{I})$ helps support probabilistic forecasting. Finally, we can obtain the forecasts $\hat{y}_{i}\in\mathbb{R}^{p^{time}}$ of the $i$-th future token. And the forecasts of the future horizon are $\hat{Y}=\text{Concat}\{\hat{y}_{i}\}\in\mathbb{R}^{F\times p^{time}}$.
@@ 4.1 Experimental Settings
Baselines. We compare Aurora with 11 well-known unimodal time series foundation models, including Sundial (Liu et al., 2025e), VisionTS (Chen et al., 2024a), ROSE (Wang et al., 2025c), Time-MoE (Shi et al., 2024b), MOIRAI (Woo et al., 2024), TTM (Ekambaram et al., 2024), TimesFM (Das et al., 2024), Timer (Liu et al., 2024e), UniTS (Gao et al., 2024), Chronos (Ansari et al., ), and Lag-Llama (Rasul et al., 2023). We also consider multiple strong end-to-end supervised models, including multimodal ones like GPT4MTS (Jia et al., 2024), TATS (Li et al., 2025d), CALF (Liu et al., 2025a), and Time-VLM (Zhong et al., ), and unimodal ones like TimeXer (Wang et al., 2024b), PatchTST (Nie et al., 2023), iTransformer (Liu et al., 2024d) TSDiff (Kollovieh et al., 2023), CSDI (Tashiro et al., 2021), TimeGrad (Rasul et al., 2021a), and GRU MAF (Rasul et al., 2021b). The detailed information is provided in Appendix A.
@@ A.1 Cross-domain Multimodal Time Series Corpus
Textual Descriptions. Since multimodal time series data is scarce in the real world, current works (Liu et al., 2024b; Kong et al., 2025; Xie et al., 2024) often construct corresponding textual information based on human experiences and Large Language Models, which is proven effective in training models. Following their paradigm, we provide raw time series data in Figure 7 with domain descriptions, encouraging GPT4 (Achiam et al., 2023) to heuristically generate textual descriptions of sample-wise time series, thus obtaining high-quality multimodal time-series data from simulation. Specifically, after a GPT4 agent generates the textual descriptions, we first coarsely check the quality with another GPT4 agent. If the quality is low, the process will be reset. After a batch of textual descriptions are generated, we randomly sample from them and check the quality manually, then determine whether to regenerate this batch of data and tune the prompts. As shown in Figure 8, here’s some samples of our prompts for GTP4 and the generated texts:
@@ A.2 Baselines
For zero-shot forecasting, we compare Aurora against 11 advanced foundation models: Sundial (Liu et al., 2025e), VisionTS (Chen et al., 2024a), ROSE (Wang et al., 2025c), Timer (Liu et al., 2024e), MOIRAI (Woo et al., 2024), TTM (Olivares et al., 2023), Chronos (Ansari et al., ), TimesFM (Das et al., 2024), Time-MoE (Shi et al., 2024b), UniTS (Gao et al., 2024), and Lag-Llama (Rasul et al., 2023). We also compare Aurora with total multiple End-to-end supervised models: TimeXer (Wang et al., 2024b), PatchTST (Nie et al., 2023), iTransformer (Liu et al., 2024d), TimesNet (Wu et al., 2023), GPT4MTS (Jia et al., 2024), TATS (Li et al., 2025d), CALF (Liu et al., 2025a), Time-VLM (Zhong et al., ), TSDiff (Kollovieh et al., 2023), TimeGrad (Rasul et al., 2021a), CSDI (Tashiro et al., 2021), and GRU MAF (Rasul et al., 2021b). The corresponding codebases and implementation details are summarized in Table 6.
@@ Appendix E Related Works
Time Series Analysis assumes a position of utmost significance within a wide array of fields, including the economy (Qiu et al., 2025b; a; Liu et al., 2025a), transportation (Wu et al., 2025f; 2024), health (Lu et al., 2023; 2024a; 2024b), weather (Li et al., 2025a; Yang et al., 2024; Zhou et al., 2025), and energy (Sun et al., 2025a). It encompasses numerous crucial tasks, such as forecasting (Li et al., 2025b; Dai et al., 2024c), anomaly detection (Wang et al., 2025a; Qiu et al., 2025c; Wu et al., 2025e), classification (Liu et al., 2024d), imputation (Wu et al., 2023; 2025b), and others (Qiu et al., 2025e; Liu et al., 2026; 2025c). Among these tasks, Time Series Forecasting is the one most widely applied in real-world situations.
Time series forecasting (TSF) entails the prediction of future observations based on historical ones. Research findings have indicated that features learned through certain methods may exhibit superior performance compared to human-designed features (Hu et al., 2025a; Dai et al., 2024c; Liu et al., 2025d; Sun et al., 2025b; a; Niu et al., 2025). By harnessing the representation learning capabilities of deep neural networks (DNNs), a multitude of deep-learning based approaches have emerged. For instance, TimesNet (Wu et al., 2023) and SegRNN (Lin et al., 2023) model time series as vector sequences, employing CNNs or RNNs to capture temporal dependencies. Moreover, Transformer architectures, including Informer (Zhou et al., 2021), Dsformer (Yu et al., 2023), TimeFilter (Hu et al., 2025b), TimeBridge (Liu et al., 2025b), PDF (Dai et al., 2024b), Triformer (Cirstea et al., 2022), PatchTST (Nie et al., 2023), ROSE (Wang et al., 2025c), LightGTS (Wang et al., 2025b), and FLAME (Wu et al., 2025a), are capable of more accurately capturing the complex relationships between time points, thereby significantly enhancing forecasting performance. MLP-based methods, such as DUET (Qiu et al., 2025d), $K^{2}$VAE (Wu et al., 2025d), SRSNet (Wu et al., 2025c), AMD (Hu et al., 2025a), SparseTSF (Lin et al., 2024a), CycleNet (Lin et al., 2024b), NLinear (Zeng et al., 2023), and DLinear (Zeng et al., 2023), adopt relatively simpler architectures with fewer parameters yet still manage to achieve highly competitive forecasting accuracy.

### v5

@@ Abstract
Cross-domain generalization is very important in Time Series Forecasting because similar historical information may lead to distinct future trends due to the domain-specific characteristics. Recent works focus on building unimodal time series foundation models and end-to-end multimodal supervised models. Since domain-specific knowledge is often contained in modalities like texts, the former lacks the explicit utilization of them, thus hindering the performance. The latter is tailored for end-to-end scenarios and does not support zero-shot inference for cross-domain scenarios. In this work, we introduce Aurora, the first Multimodal Time Series Foundation Model, which supports multimodal inputs and zero-shot inference. Pretrained on Cross-domain Multimodal Time Series Corpus, Aurora can adaptively extract and focus on key domain knowledge contained in corresponding text or image modalities, thus possessing strong cross-domain generalization capability. Through tokenization, encoding, and distillation, Aurora can extract multimodal domain knowledge as guidance and then utilizes a Modality-Guided Multi-head Self-Attention to inject them into the modeling of temporal representations. In the decoding phase, the multimodal representations are used to generate the conditions and prototypes of future tokens, contributing to a novel Prototype-Guided Flow Matching for generative probabilistic forecasting. Comprehensive experiments on 5 well-recognized benchmarks, including TimeMMD, TSFM-Bench, ProbTS, TFB, and EPF, demonstrate the consistent state-of-the-art performance of Aurora on both unimodal and multimodal scenarios.
@@ 1 Introduction
Time series forecasting has gained sustained attention for decades of years due to its significant values in multiple domains, including economy, transportation, meteorology, and public health. In recent years, the key pivot comes with the surge of deep learning, which brings the boom of merticulously-designed deep forecasting models (Cirstea et al., 2022; Nie et al., 2023; Qiu et al., 2025e; Wu et al., 2025d). Through learning the inherent dynamics within the raw data, deep learning models can outperform classic statistical methods (Box & Pierce, 1970; Mei et al., 2014) and obey the scaling law (Shi et al., 2024a; Yao et al., ). Due to the success, it also brings the most commonly-used forecasting paradigm, which utilizes the past information to infer how the series goes in the coming horizon. Although this paradigm contributes to impressive performance under the domain-specific scenarios, its effectiveness is uncertain when facing cross-domain inference, where similar historical information may lead to different futures due to domain differences.
@@
2) utilizing cross-modality information in training end-to-end multimodal supervised models, which effectively integrates domain knowledge in forecasting. For time series foundation models, the cross-domain generalization capabilities mainly come from the sensitivity to subtle differences in historical information from different domains. Some of them (Shi et al., 2024b; Liu et al., 2025e) are pretrained on trillion-scale corpus with hevay backbones, thus possessing certain cross-domain adaption capabilities. Others (Wang et al., 2025f; e; Wu et al., 2025a) have specific structures, which excels at capturing cross-domain features. However, their capabilites come from single time modality and lack explict domain knowledge guidance, thus hindering the performance. For end-to-end multimodal supervised models (Jin et al., 2024; Liu et al., 2025b), though they consider the multimodal knowledge to enhance the domain-specific forecasting, they lack the ability to support zero-shot forecasting in cross-domain scenarios. In our view, the aurora of next-generation time series foundation model lies in pretraining a cross-modality model on cross-domain time series corpus, which can utilize the domain knowledge within modalities and serve as a versatile out-of-the-box forecaster in complex scenarios.
See Figure 1, we propose Aurora, which pioneers the exploration of multimodal time series foundation model. Specifically, we pretrain Aurora on Cross-Domain Multimodal Time Series Corpus, with time series data and sample-wise domain-specific text descriptions. Since previous works (Chen et al., 2024a; Yu et al., 2025c) point out the endogenous images of time series contain additional geometric information, we also consider them into cross-modality learning. Considering the model architecture, Aurora adopts a novel cross-modality Encoder. Taking pretrained Bert (Devlin et al., 2019) and ViT (Liu et al., 2021) as modality encoders, Aurora then adopts token distillation to extract the key information in different modalities. To effectively model the cross-modality interaction, we propose a novel Modality-Guided Self-Attention Mechanism to utilize the external domain knowledge to adjust the attention of internal information within the time series data to obtain temporal features, and then fuse them with text and image features.
@@ 2.1 Time Series Forecasting
Time Series Forecasting is vital in decision-making and has fascinated people for decades of years, which facilitates the emergence of a series of works. In recent years, deep-learning models are widely studied, among them, Autoformer (Wu et al., 2021), Triformer (Cirstea et al., 2022), TimesNet (Wu et al., 2023), Pathformer (Chen et al., 2024b), PatchTST (Nie et al., 2023), Dlinear (Zeng et al., 2023), FiTS (Xu et al., 2024), SparseTSF (Lin et al., 2024a), PDF (Dai et al., 2024a), DUET (Qiu et al., 2025e), and TimeMixer++ (Wang et al., ), continuously advancing the state-of-the-arts. However, though they possess the capabilities to extract the inherent dynamics in raw time series data, they only adapt to unimodal end-to-end forecasting scenarios, and often fall short in multimodal forecasting scenarios where the domain knowledge is widely contained in the text modality.
Recently, some works are proposed to explore the multimodal end-to-end supervised models. In summary, they utilize Large Language Models’ strong reasoning capabilities to integrate textual domain knowledge to prompt temporal modeling. Among them, Unitime (Liu et al., 2024c), Time-LLM (Jin et al., 2024) and CALF (Liu et al., 2025b) utilize the endogenous textual descriptions as prompts, GTP4MTS (Jia et al., 2024), TATS (Li et al., 2025d) and TimeMMD (Liu et al., 2024b) supports exogenous textual domain knowledge. However, they do not possess generalization capabilities in zero-shot scenarios.
@@ 2.2 Time Series Foundation Models
To support cross-domain generalization, unimodal Time Series Foundation Models are widely studied. The majority of them adopt Tranformer-based architectures, which are pretrained on time series corpus of billion- or trillion- scale to obtain the strong generalization capabilities. Among them, Sundial (Liu et al., 2025e), VisionTS (Chen et al., 2024a), ROSE (Wang et al., 2025f), Time-MoE (Shi et al., 2024b), MOIRAI (Woo et al., 2024), TTM (Ekambaram et al., 2024), Chronos (Ansari et al., ), UniTS (Gao et al., 2024), Timer (Liu et al., 2024e), and TimesFM (Das et al., 2024) demonstrate strong zero-shot forecasting performance on unimodal tasks, even outperforming those full-shot supervised models in many cases. Considering the forecasting paradigm, Sundial, MOIRAI, Chronos, and Lag-Llama (Rasul et al., 2023) also support probabilistic forecasting, which provides additional robustness and versatility for decision-making. Despite their endeavors to enhance cross-domain generalization capabilities, when historical series exhibit similarities, the forecasts they generate remain static. This lack of adaptability renders them unable to accommodate the diverse and changing real-world domains.
@@ 3.1.3 Multimodal Alignment
where VAttn and TAttn are unnormalized attention scores, separately denoting the correlations between time modality and image or text modality. $\text{Corr}\in\mathbb{R}^{n^{time}\times n^{time}}$ denotes the inherent temporal correlations bridged through the cross-modality correlations. We also introduce $W\in\mathbb{R}^{K^{image}\times K^{text}}$ as a learnable metric (Qiu et al., 2025e) to further tune the semantic distances. This process can help bridge the correlations between time series tokens through multimodal information, capable of leveraging domain knowledege to guide the modeling of temporal dynamics. We then inject Corr into the temporal encoding process:
@@ 3.2.2 Prototype-Guided Flow Matching
where $\mathcal{D}$ denotes the weights of prototypes, we then generate the new prototypes through: $\tilde{\mathcal{P}}=\mathcal{D}\cdot\mathcal{P}\in\mathbb{R}^{F\times p^{time}}$, where the generated prototype $\tilde{\mathcal{P}}$ contains the approximate future periodicities and trends. As shown in Figure 3, the motivation of Flow Matching is to fit the velocity field between the initial prototype $y_{i}^{(0)}=\tilde{\mathcal{P}}_{i}+\epsilon_{i}$ and the target horizon $y_{i}^{(1)}=\mathrm{y}_{i}$, where $\mathrm{y}_{i}\in\mathbb{R}^{p^{time}}$ is the groundtruth of the $i$-th future token, and $\epsilon_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is used to increase the diversity during training. We design the Flow-Matching Network with an MLP structure and utilize the AdaLN (Peebles & Xie, 2023) to integrate the multimodal conditions $h_{i}=\mathrm{X}^{cond}_{i}$. We adopt the conditional optimal-transport path, which is energy-optimal and contributes to a uniform velocity field. And the function of Flow-Matching Network $v^{\theta}_{t}$ is to predict the velocity based on the current position $y_{i}^{(t)}$ and condition $h_{i}$. To achieve this, the token-wise optimization objective $\mathcal{L}$ is designed as:
@@
In the inference phase, the sampling process is a discretized integration process–see Algorithm 3.2.2. The Gaussian noise $\epsilon_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ helps support probabilistic forecasting. Finally, we can obtain the forecasts $\hat{y}_{i}\in\mathbb{R}^{p^{time}}$ of the $i$-th future token. And the forecasts of the future horizon are $\hat{Y}=\text{Concat}\{\hat{y}_{i}\}\in\mathbb{R}^{F\times p^{time}}$.
@@ 4.1 Experimental Settings
Baselines. We compare Aurora with 11 well-known unimodal time series foundation models, including Sundial (Liu et al., 2025e), VisionTS (Chen et al., 2024a), ROSE (Wang et al., 2025f), Time-MoE (Shi et al., 2024b), MOIRAI (Woo et al., 2024), TTM (Ekambaram et al., 2024), TimesFM (Das et al., 2024), Timer (Liu et al., 2024e), UniTS (Gao et al., 2024), Chronos (Ansari et al., ), and Lag-Llama (Rasul et al., 2023). We also consider multiple strong end-to-end supervised models, including multimodal ones like GPT4MTS (Jia et al., 2024), TATS (Li et al., 2025d), CALF (Liu et al., 2025b), and Time-VLM (Zhong et al., ), and unimodal ones like TimeXer (Wang et al., 2024b), PatchTST (Nie et al., 2023), iTransformer (Liu et al., 2024d) TSDiff (Kollovieh et al., 2023), CSDI (Tashiro et al., 2021), TimeGrad (Rasul et al., 2021a), and GRU MAF (Rasul et al., 2021b). The detailed information is provided in Appendix A.
@@ A.1 Cross-domain Multimodal Time Series Corpus
Textual Descriptions. Since multimodal time series data is scarce in the real world, current works (Liu et al., 2024b; Kong et al., 2025; Xie et al., 2024) often construct corresponding textual information based on human experiences and Large Language Models, which is proven effective in training models. Following their paradigm, we provide raw time series data in Figure 7 with domain descriptions, encouraging GPT-4 (Achiam et al., 2023) to heuristically generate textual descriptions of sample-wise time series, thus obtaining high-quality multimodal time-series data from simulation. Specifically, after a GPT-4 agent generates the textual descriptions, we first coarsely check the quality with another GPT-4 agent. If the quality is low, the process will be reset. After a batch of textual descriptions are generated, we randomly sample from them and check the quality manually, then determine whether to regenerate this batch of data and tune the prompts. As shown in Figure 8, here’s some samples of our prompts for GTP4 and the generated texts:
@@ A.2 Baselines
For zero-shot forecasting, we compare Aurora against 11 advanced foundation models: Sundial (Liu et al., 2025e), VisionTS (Chen et al., 2024a), ROSE (Wang et al., 2025f), Timer (Liu et al., 2024e), MOIRAI (Woo et al., 2024), TTM (Olivares et al., 2023), Chronos (Ansari et al., ), TimesFM (Das et al., 2024), Time-MoE (Shi et al., 2024b), UniTS (Gao et al., 2024), and Lag-Llama (Rasul et al., 2023). We also compare Aurora with total multiple End-to-end supervised models: TimeXer (Wang et al., 2024b), PatchTST (Nie et al., 2023), iTransformer (Liu et al., 2024d), TimesNet (Wu et al., 2023), GPT4MTS (Jia et al., 2024), TATS (Li et al., 2025d), CALF (Liu et al., 2025b), Time-VLM (Zhong et al., ), TSDiff (Kollovieh et al., 2023), TimeGrad (Rasul et al., 2021a), CSDI (Tashiro et al., 2021), and GRU MAF (Rasul et al., 2021b). The corresponding codebases and implementation details are summarized in Table 6.
@@ Appendix E Related Works
Time Series Analysis assumes a position of utmost significance within a wide array of fields, including the economy (Qiu et al., 2025b; a; Liu et al., 2025b), transportation (Wu et al., 2025f; 2024; Miao et al., 2025; Liu et al., 2025a), health (Lu et al., 2023; Miao et al., 2024a; Lu et al., 2024; Wang et al., 2026a), weather (Li et al., 2025a; Miao et al., 2024b; Zhou et al., 2025), and energy (Sun et al., 2025a). It encompasses numerous crucial tasks, such as forecasting (Li et al., 2025b; Dai et al., 2024c; Shang et al., 2026; 2024), anomaly detection (Wang et al., 2025d; Qiu et al., 2025c; Wu et al., 2025e), classification (Liu et al., 2024d), imputation (Wu et al., 2023; 2025b; Wang et al., 2025a; Shang & Chen, 2024), irregular forecasting (Liu et al., 2026b; a; Yu et al., 2025a), motion forecasting (Ma et al., 2025a; 2024a; 2024b; c; b) and others (Qiu et al., 2025f; Wang et al., 2023; Yu et al., 2025b; Cheng et al., 2026). Among these tasks, Time Series Forecasting is the one most widely applied in real-world situations.
Time series forecasting (TSF) entails the prediction of future observations based on historical ones. Research findings have indicated that features learned through certain methods may exhibit superior performance compared to human-designed features (Liu et al., 2025d; Sun et al., 2025b; a; Niu et al., 2025; Chen et al., 2023). By harnessing the representation learning capabilities of deep neural networks (DNNs), a multitude of deep-learning based approaches have emerged. For instance, TimesNet (Wu et al., 2023) and SegRNN (Lin et al., 2023) model time series as vector sequences, employing CNNs or RNNs to capture temporal dependencies. Moreover, Transformer architectures, including Informer (Zhou et al., 2021), Dsformer (Yu et al., 2023), TimeFilter (Hu et al., 2025b), TimeBridge (Liu et al., 2025c), PDF (Dai et al., 2024b), Triformer (Cirstea et al., 2022), PatchTST (Nie et al., 2023), ROSE (Wang et al., 2025f), LightGTS (Wang et al., 2025e), and FLAME (Wu et al., 2025a), are capable of more accurately capturing the complex relationships between time points, thereby significantly enhancing forecasting performance. MLP-based methods, such as DUET (Qiu et al., 2025e), $K^{2}$VAE (Wu et al., 2025d), SRSNet (Wu et al., 2025c), AMD (Hu et al., 2025a), SparseTSF (Lin et al., 2024a), CycleNet (Lin et al., 2024b), NLinear (Zeng et al., 2023), and DLinear (Zeng et al., 2023), adopt relatively simpler architectures with fewer parameters yet still manage to achieve highly competitive forecasting accuracy. Other approaches model the inductive biases within temporal data more effectively from the perspective of loss functions, such as DBLoss (Qiu et al., 2025d), Time-O1 (Wang et al., 2025c), FreDF (Wang et al., 2025b), and DsitDF (Wang et al., 2026b).